# Towards Open-World Retrieval-Augmented Generation on Knowledge Graph: A Multi-Agent Collaboration Framework

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated strong capabilities in language understanding and reasoning. However, their dependence on static training corpora makes them prone to factual errors and knowledge gaps. Retrieval-Augmented Generation (RAG) addresses this limitation by incorporating external knowledge sources, especially structured Knowledge Graphs (KGs), which provide explicit semantics and efficient retrieval. Existing KG-based RAG approaches, however, generally assume that anchor entities are accessible to initiate graph traversal, which limits their robustness in open world settings where accurate linking between the query and the entity is unreliable. To overcome this limitation, we propose AnchorRAG, a novel multi-agent collaboration framework for open-world RAG without the predefined anchor entities. Specifically, a predictor agent dynamically identifies candidate anchor entities by aligning user query terms with KG nodes and initializes independent retriever agents to conduct parallel multi-hop explorations from each candidate. Then a supervisor agent formulates the iterative retrieval strategy for these retriever agents and synthesizes the resulting knowledge paths to generate the final answer. This multi-agent collaboration framework improves retrieval robustness and mitigates the impact of ambiguous or erroneous anchors. Extensive experiments on four public benchmarks demonstrate that AnchorRAG significantly outperforms existing baselines and establishes new state-of-the-art results on the real-world question answering tasks.
1 Introduction
@@
Large Language Models (LLMs) (Achiam et al. 2023; Yang et al. 2025; Liu et al. 2024) are typically defined as deep learning models with a massive number of parameters, trained on large-scale corpora in a self-supervised manner. Their internal parameterization allows for an implicit representation of external knowledge. During the inference stage, the Chain-of-Thought (CoT) (Wei et al. 2022; Trivedi et al. 2022; Zhou et al. 2022) method guides the models to “think step-by-step” through carefully designed prompts. It helps the models better handle logically complex or multi-step reasoning tasks, achieving a better performance in natural language processing tasks such as question answering (Gu et al. 2021). Despite these advances, LLMs are fundamentally limited by the incompleteness and inaccuracies in their static training corpora. This often results in factual hallucinations (Huang et al. 2025), especially in knowledge-intensive tasks, significantly hindering their reliability and deployment in real-world applications.
Figure 1: Illustration of the difference between the closed-world and the open-world setting in RAG.
@@
To address these limitations, recent studies (Jeong et al. 2024; Xia et al. 2025) have investigated augmenting LLMs with external knowledge sources to improve factual consistency and reasoning. While fine-tuning (Hu et al. 2022; Li and Liang 2021; Li et al. 2023b), LLM with the external knowledge is computationally expensive and inflexible, a promising solution is Retrieval-Augmented Generation (RAG) (Lewis et al. 2020; Jiang et al. 2023b; Sarthi et al. 2024), which combines real-time retrieval from external sources with generative modeling. Among exiting methods, RAG on the Knowledge Graph (KG) (Jiang et al. 2023a; Ma et al. 2025; He et al. 2024; Chen et al. 2024; Guo et al. 2024) has become a research hotspot in the field, due to the fact that KGs can provide structured knowledge and explicit retrieval paths which benefit to enhancing the reasoning capabilities of LLMs. Typically, these KG-based RAG methods (Wang et al. 2025; Sun et al. 2023; Ma et al. 2024) can be divided in two steps: (1) identifying a query-aware anchor entity to initiate retrieval; (2) performing iterative retrieval to construct the reasoning paths. Through above steps, these methods can effectively exploit the external knowledge related to the query and improve the performance in the question answering task.
@@
However, existing KG-based RAG methods almost follow the closed-world setting where the anchor entity is accessible and exists in the given KG. As shown in the Fig. 1 (a), the anchor entity “Mitt Romney” is pre-defined for the question “What university did Romney graduated from?”. The RAG methods can directly locate the target entity for the following retrieval. Actually, user questions usually conform to the open-world setting where the anchor entity is unavailable. In that case, existing methods extract the keywords in the query and identify the candidate anchor entity by semantic matching. Due to name abbreviations and aliases, these methods usually suffer from the issue caused by imprecise or partial matching as shown in Fig. 1(b). Then the challenge of the open-world RAG lies in how to accurately identify anchor entities to retrieve the relevant information.
@@
  • Methodologies. For anchor identifying, we design an entity grounding strategy to locate the candidate anchors by entity name description and structural neighborhood. For knowledge retrieval, we propose a novel retrieval method with rough pruning and fine filtering, which can capture the resultful knowledge paths to boost the LLM reasoning.
@@
  • Experimental Findings. Extensive experiments demonstrate that AnchorRAG consistently outperforms existing baselines on four public QA datasets, establishing state-of-the-art results on the real-world question answering tasks.
@@
Figure 2: An overview of the AnchorRAG framework, including the collaborative workflows of multi-agent exploration for question answering.
2 Related Work
While advanced prompting methods like Chain-of-Thought (Wei et al. 2022) enhance LLM reasoning, their reliance on static internal knowledge can lead to factual errors and hallucinations that accumulate over multi-step inference. Retrieval-Augmented Generation (RAG) (Lewis et al. 2020; Semnani et al. 2023; He, Zhang, and Roth 2022) mitigates this by leveraging external knowledge, significantly improving performance on knowledge-intensive tasks like Knowledge Graph Question Answering (KGQA)  (Yih et al. 2016), where the knowledge graphs (KGs) (Bollacker et al. 2008; Auer et al. 2007; Vrandečić and Krötzsch 2014) provide large-scale structured facts and rich semantic relationships. These KG-based RAG methods can be divided into two main groups: KG-augmented Fine-tuning and KG-augmented In-context Learning.
2.1 KG-augmented Fine-tuning
KG-augmented fine-tuning directly integrates structured knowledge into open-source LLMs during training, operating at the entity, path, or subgraph levels (Zhang et al. 2025; Pan et al. 2024). At the entity level, methods like ChatDoctor (Li et al. 2023a) and PMC-LLaMA (Wu et al. 2024) employ instruction tuning (Wang et al. 2022) to enhance domain-specific knowledge understanding before fine-tuning on downstream tasks that contain vast entities and text descriptions. For path-level fine-tuning, RoG (Luo et al. 2023b) trains the LLM to generate valid relation paths, thereby aligning the reasoning process with the graph’s actual structure. The sources of the knowledge paths is diverse, ranging from the direct routes between question-related entities to complex logical rules extracted by graph retrieval models (Mavromatis and Karypis 2024) or heuristic strategies (Tan et al. 2024). To capture more comprehensive structural information, subgraph-level methods fine-tune LLMs on the entire local neighborhood, compressing them into embeddings (He et al. 2024) or transforming them into text sequences (Luo et al. 2023a).
2.2 KG-augmented In-context Learning
For state-of-the-art closed-source LLMs where fine-tuning is infeasible (Zhang et al. 2025), KG-augmented In-context Learning (ICL) is the primary strategy (Huang et al. 2023). Early methods like ToG (Sun et al. 2023) treat the LLM as an agent that executes sequential reasoning along the retrieved relational path. Given that this methods are prone to error accumulation and inefficient search, subsequent research has developed more advanced reasoning frameworks. To enhance reliability, Plan-on-Graph (Chen et al. 2024) introduced adaptive planning and self-correction mechanism, while Debate-on-Graph (Ma et al. 2025) proposed multi-path argumentation. Besides, recent methods, e.g., Fast Think-on-Graph  (Liang and Gu 2025), focused on optimizing search efficiency, they accelerates the reasoning process and improves accuracy by enabling LLMs to “think community-by-community” within the knowledge graph. To address knowledge graph incompleteness, ReKnoS (Wang et al. 2025) simplify reasoning by utilizing “super-relations”, while others, e.g., Generate-on-Graph (Xu et al. 2024), dynamically generate facts when external knowledge is unavailable. Nevertheless, almost of the existing methods are implemented for the closed-world question answering, which limits their performance in the real-world settings where accurate linking between the query and the entity is unreliable.
3 Method
3.1 Task Formulation
The fundamental unit of data in a knowledge graph $\mathcal{G}$ is the triples, i.e., $\mathcal{G}=\{(h,r,t)\mid h,t\in\mathcal{E},r\in\mathcal{R}\}$, where $\mathcal{E}$ and $\mathcal{R}$ respectively represent the set of entities and relations.The goal of Knowledge Graph Question Answering (KGQA) is to answer natural language questions by leveraging the structured triples within a knowledge graph. An unavoidable challenge for KGQA is how to design a synergy between the semantic understanding of LLMs and the structured facts of a knowledge graph, enabling both precise knowledge retrieval and powerful reasoning for the QA task.
3.2 Overview
As shown in the Fig. 2, our framework follows a pipeline-based MAC framework: anchor identifying by the predictor agent, knowledge searching by the retriever agents and question answering by the supervisor agent. Specifically, the proposed predictor agent first extracts the keywords from the given question and applies a semantic match model to obtain the candidate entities by their name description and structural neighborhood. Then, multiple retriever agents initiate parallel graph traversal for each candidate, performing iterative retrieval and obtaining the pivotal knowledge for the following reasoning. Furthermore, at each retrieval round, a supervisor agent aggregates the retrieved evidence and determines whether an answer can be generated and the retrieval process should be continued.
3.3 Anchor Entity Identifying
@@
In our framework, the process of accurately identifying anchor entities from natural language questions is handled by the predictor agent. This process consists of the following two steps:
Candidate Entity Generation:
The predictor agent is associated with a LLM to recognize and extract keywords from the input question $q$, forming a keyword set $\mathcal{D}=\{d_{1},d_{2},\ldots,d_{n}\}$. For example, given the question $q$: ”What country bordering France contains an airport that serves Nijmegen?”, the predictor locates the keyword set $\mathcal{D}$, i.e., $\mathcal{D}=\{\text{France},\text{Nijmegen}\}$. In addition, considering there may exist certain spelling errors in the open-world questions, then the predictor will proofread the given question to filter out irrelevant distractors and correct potential errors, ensuring that the extracted keywords are clean and meaningful. Subsequently, it adopts a semantic match between these keywords and the entities set $\mathcal{E}$ of the external KG, generating an initial list of candidate anchors $\tilde{\mathcal{E}}$ through in-context learning. The prompt and in-context examples can be found in the Appendix.
Relation-Aware Entity Grounding:
@@
The initial candidate set $\tilde{\mathcal{E}}$ often contains numerous entities with similar names that are irrelevant to the question’s context. Taking all of them into account will increase the overhead of subsequent retrieval and inevitably introduce noise. In this section, we argue that an entity is a more probable anchor if the semantic implied in its neighborhood is aligned with the given question. Therefore, the predictor agent presents a relation-aware grounding strategy to precisely identify the most relevant candidate entities. For each candidate entity $e\in\tilde{\mathcal{E}}$, the agent gathers its one-hop neighboring relations and obtains the set $\mathcal{R}(e)$. Then it invokes a pre-trained model SBERT (Reimers and Gurevych 2019) to encode the neighboring relations and the question, obtaining the relevance score $s(r_{i},q)$ between each relation $r_{i}\in\mathcal{R}(e)$ and the given question $q$ as follows:
$s(r_{i},q)=\left\langle\mathrm{SBERT}(r_{i}),\mathrm{SBERT}(q)\right\rangle,$ (1)
where $\left\langle\,\right\rangle$ denotes the inner product. Based on the above equation, we define the final relevance score of the candidate entity by aggregating the semantic information of its neighboring relations as follows:
$\mathrm{Score}(e,q)=\frac{1}{k}\sum_{r\in\mathcal{N}(e)}s(r,q),$ (2)

### v2

@@ Abstract.
Large Language Models (LLMs) have demonstrated strong capabilities in web search and reasoning. However, their dependence on static training corpora makes them prone to factual errors and knowledge gaps. Retrieval-Augmented Generation (RAG) addresses this limitation by incorporating external knowledge sources, especially structured Knowledge Graphs (KGs), which provide explicit semantics and efficient retrieval. Existing KG-based RAG approaches, however, generally assume that anchor entities are accessible to initiate graph traversal, which limits their robustness in open-world settings where accurate linking between the user query and the KG entity is unreliable. To overcome this limitation, we propose AnchorRAG, a novel multi-agent collaboration framework for open-world RAG without the predefined anchor entities. Specifically, a predictor agent dynamically identifies candidate anchor entities by aligning user query terms with KG nodes and initializes independent retriever agents to conduct parallel multi-hop explorations from each candidate. Then a supervisor agent formulates the iterative retrieval strategy for these retriever agents and synthesizes the resulting knowledge paths to generate the final answer. This multi-agent collaboration framework improves retrieval robustness and mitigates the impact of ambiguous or erroneous anchors. Extensive experiments on four public benchmarks demonstrate that AnchorRAG significantly outperforms existing baselines and establishes new state-of-the-art results on the real-world reasoning tasks. The datasets along with our code are available at https://github.com/Jayson3831/AnchorRAG.
Retrieval-Augmented Generation; Knowledge Graph; Large Language Model; Agentic search
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emirates.booktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesisbn: 979-8-4007-2307-0/2026/04doi: 10.1145/3774904.3792389ccs: Information systems Language models
Resource Availability:
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.18310915.
1. Introduction
@@
Large Language Models (LLMs) (Achiam et al., 2023; Yang et al., 2025) are typically defined as deep learning models with a massive number of parameters, trained on large-scale corpora in a self-supervised manner. Their internal parameterization allows for an implicit representation of external knowledge. During the inference stage, the chain-of-Thought (CoT) (Wei et al., 2022) method guides the models to “think step-by-step” through carefully designed prompts. It helps the models better handle logically complex or multi-step reasoning tasks, achieving a better performance in natural language processing tasks such as web search (Li et al., 2025). Despite these advances, LLMs are fundamentally limited by the incompleteness and inaccuracies in their static training corpora. This often results in factual hallucinations (Huang et al., 2025), especially in knowledge-intensive tasks, significantly hindering their reliability and deployment in real-world applications.
@@
To address these limitations, recent studies (Jeong et al., 2024; Xia et al., 2025) have investigated augmenting LLMs with external knowledge sources to improve factual consistency and reasoning. While fine-tuning (Hu et al., 2022) LLM with the external knowledge is computationally expensive and inflexible, a promising solution is Retrieval-Augmented Generation (RAG) (Lewis et al., 2020), which combines real-time retrieval from external sources with generative modeling. To improve the retrieval efficiency of the external sources, some researchers apply the knowledge graph (KG) to represent the external information as the structured triples, i.e., (entity, relation, entity). Then the core issue of RAG on the knowledge graph is to locate the relevant triples for enhancing the reasoning capability.
Among existing solutions, iterative retrieval methods (Sun et al., 2023; Ma et al., 2025; Chen et al., 2024; Xu et al., 2024; Wang et al., 2025) has become a research hotspot in the field, due to the fact that they can handle the complex user queries by dynamically constructing and refining explainable reasoning paths through multi-round “retrieval-generation”. Typically, these RAG methods can be divided in two steps: (1) identifying a query-aware anchor entity to initiate retrieval; (2) performing iterative retrieval to construct the reasoning paths. Through the above steps, these methods can effectively exploit the external knowledge related to the query and improve the performance in the downstream reasoning task.
Figure 1. Illustration of the difference between the closed-world and the open-world setting in RAG.
@@
However, existing KG-based RAG methods almost follow the closed-world setting where the anchor entity is accessible and exists in the given KG. As shown in the Figure 1 (a), the anchor entity “Mitt Romney” is pre-defined for the question “What university did Romney graduated from?”. The RAG methods can directly locate the target entity for the following retrieval. Actually, user questions usually conform to the open-world setting where the anchor entity is unavailable. In that case, existing methods usually borrow from entity linking models and extract the keywords in the query to identify the candidate anchor entity by semantic matching. Due to name abbreviations and aliases, these methods usually suffer from the issue caused by imprecise or partial matching as shown in Figure 1(b). Then the challenge of the open-world RAG lies in how to accurately identify anchor entities to retrieve the relevant information.
@@
  • Methodologies. For anchor identifying, we design an entity grounding strategy to locate the candidate anchors by entity name description and structural neighborhood. For knowledge retrieval, we propose a novel retrieval method with candidate ranking and evidence validation, which can capture the resultful knowledge paths to boost the LLM reasoning.
@@
  • Experimental Findings. Extensive experiments demonstrate that AnchorRAG consistently outperforms existing baselines on four public QA datasets, establishing state-of-the-art results on the real-world reasoning tasks.
2. Related Work
Knowledge graphs(Bollacker et al., 2008; Auer et al., 2007) provide essential supplementation to the static knowledge of LLMs(Gao et al., 2023), effectively alleviating factual inaccuracies and hallucination issues during complex reasoning(Yih et al., 2016) tasks. Consequently, the efficient retrieval of factual information from KGs for LLMs has become a critical topic of research, which can be divided into three main groups: Semantic Parsing, Single-Pass Retrieval and Iterative Retrieval methods.
2.1. Semantic Parsing Methods
Semantic parsing methods aim to convert natural language questions into executable logical forms(Lan et al., 2022), such as S-expressions or SPARQL queries. The results from these queries are then used as external knowledge to enhance the LLM’s reasoning capabilities. Researchers have explored various strategies for generating and evaluating these logical forms. FlexKBQA(Li et al., 2024) first samples a multitude of S-expressions and employs a LLM to convert them into natural language questions, creating program-question pairs which are then used for training lightweight models. Another typical method RNG-KBQA(Ye et al., 2022) resorts to a ranking model to select the most relevant logical forms based on a set of candidates. Additionally, ChatKBQA(Luo et al., 2023) introduces a “Generate-then-Retrieve” framework, which first uses a fine-tuned LLM to generate a logical form skeleton and then populates it with specific details via an unsupervised retrieval module. It is important to note that if the generated logical forms are inaccurate, the entire reasoning pipeline will be confused, leading to incorrect final answers. This may limit the reasoning ability of semantic parsing methods in real-world scenarios, where accurate logical forms are hardly to generate.
2.2. Single-Pass Retrieval Methods
Single-Pass retrieval methods enable the acquisition of all pertinent information, including entities, relations, or subgraphs, from a knowledge graph through a single retrieval operation(Peng et al., 2024; Han et al., 2025). These methods can effectively reduce the computational costs and reasoning latency. HippoRAG (gutiérrez2025hippo) attempts to reduce the complexity and resources of retrieval process by constructing a schema-less KG, executing single-step retrieval guided by a personalized PageRank algorithm. Similarly, G-Retriever(He et al., 2024) formulates subgraph retrieval as a PCST optimization problem, creating a synergistic effect between the efficient retrieval and semantic prompting. Along this line, KnowGPT(Zhang et al., 2024) leverages reinforcement learning to extract entire reasoning chains in a single pass. Furthermore, KG-GPT(Kim et al., 2023) generates a complete evidence graph in a single instance through sentence segmentation and graph retrieval. To sum up, through above methods simplify reasoning workflows and resources, they usually face inherent limitations in tackling complex multi-hop reasoning, where an initial retrieval that fails to identify critical evidence or filter the meaningless information may easily lead to generate incorrect answers.
2.3. Iterative Retrieval Methods
Iterative retrieval frameworks are designed to tackle complex queries by dynamically constructing and refining reasoning paths through multi-round “retrieval-generation” operation(Zhang et al., 2025; Han et al., 2025). Early pioneering work, including ToG(Sun et al., 2023) and StructGPT’s IRR framework(Jiang et al., 2023), demonstrates the viability of LLMs to progressively navigate KGs and accumulate evidence. To further handle the issues caused in the highly complex reasoning, recent research resorts to more sophisticated strategies, such as multi-stage optimization in GRAG(Hu et al., 2025), dynamic knowledge completion in Generate-on-Graph(Xu et al., 2024), multi-agent deliberation in Generate-on-Graph(Ma et al., 2025), and adaptive planning and backtracking mechanisms in Plan-on-Graph(Chen et al., 2024). Despite their success, these advanced methods are fundamentally built on an idealized assumption: that entities within a query can be easily linked to their corresponding nodes in the knowledge graph. To extend to the open-world settings, some methods explore to apply entity linking strategy (Shen et al., 2015; Li et al., 2020; Vollmers et al., 2025) to identify these entities. For instance, ToG(Sun et al., 2023) and Paths-over-Graph(Tan et al., 2025) prompt LLM to automatically extract relevant entities, while AMAR(Xu et al., 2025) uses the ELQ(Li et al., 2020) model, which leverages a dual-encoder architecture to simultaneously detect mentions and link them to corresponding entities. However, these methods usually suffer from the issue caused by imprecise or partial matching, especially when the queries involving abbreviations, aliases, or long-tail entities.
3. Preliminary
Knowledge Graph is commonly defined as a directed relational graph $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{T})$, where $\mathcal{E}$ denotes the set of entities, $\mathcal{R}$ denotes the set of relations, and $\mathcal{T}\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}$ is the set of factual triples. Each triple $(h,r,t)\in\mathcal{T}$ represents a factual statement, where $h\in\mathcal{E}$ is the head entity, $r\in\mathcal{R}$ is the relation, and $t\in\mathcal{E}$ is the tail entity. For example, a triple $(\textit{Barack Obama},\textit{born\_in},\textit{Hawaii})$ encodes the fact “Barack Obama was born in Hawaii.”
Reasoning Path is defined as a sequence of linked triples in $\mathcal{G}$ that connect a source entity $e_{s}$ to a target entity $e_{t}$. Formally, a path of length $L$ can be written as:
(1) $\mathcal{P}=\begin{Bmatrix}(e_{0},r_{1},e_{1}),(e_{1},r_{2},e_{2}),\ldots,(e_{l-1},r_{l},e_{l})\end{Bmatrix},$
where $e_{0}=e_{s}$ and $e_{l}=e_{t}$. Reasoning paths capture multi-hop dependencies that are crucial for multi-hop question answering.
Reasoning Subgraph is a connected subgraph $\mathcal{G}_{q}\subseteq\mathcal{G}$ constructed during the answering process for a query $q$. Typically, $\mathcal{G}_{q}=(\mathcal{E}_{q},\mathcal{R}_{q},\mathcal{T}_{q})$ is iteratively expanded by retrieving the $k$-hop neighborhoods of frontier entities and refined to retain only the query-relevant triples. Such subgraphs serve as the evidence basis for reasoning.
Knowledge Graph Question Answering (KGQA) is the task of mapping a natural language query $q$ into one or multiple reasoning paths $\mathcal{P}$ or a reasoning subgraph $\mathcal{G}_{q}$, ultimately leading to the correct answer entity (or entities). The task requires both semantic understanding of the query and logical navigation in the KG.
Open-World KG-RAG extends the KGQA by relaxing the assumption that a set of ground-truth anchor entities $\mathcal{A}\subset\mathcal{E}$ is available for each query. Formally, given a natural language query $q$ containing a set of keywords $\mathcal{K}=\{k_{1},k_{2},\dots,k_{n}\}$, the Open-World KG-RAG is capable of autonomously learning a mapping function $\phi:\mathcal{K}\times\mathcal{G}\to\hat{\mathcal{A}}$, where $\hat{\mathcal{A}}\subset\mathcal{E}$ is the inferred candidate anchor set. Unlike the entity linking where the target entity is unique, Open-World KG-RAG is proposed for QA tasks where the target entity may not exist or may have multiple candidates.
@@ 3. Preliminary
Figure 2. An overview of the AnchorRAG framework, including the collaborative workflows of multi-agent exploration for question answering.
4. Method
In this section, we propose AnchorRAG, an iterative retrieval RAG pipeline built upon a multi-agent collaborative framework, as shown in the Figure 2. The architectural design is primarily guided by the paradigms of “divide and conquer” and “collaborative refinement”. We avoid the conventional single-threaded and error-prone “link-then-retrieve” linear paradigm. Instead, we realize the complex open-world reasoning task with three specialized sub-tasks, each operated by a distinct agent. First, the Predictor Agent is responsible for identifying anchor entities within the user queries and aligning them with the knowledge graph. Then, the Retriever Agents perform multi-hop knowledge graph retrieval and filtering, parallel to cover critical information from multiple perspectives. Finally, the Supervisor Agent supervises the above retrieval processes and generates the answers based on the information retrieved from the given KG.. The prompt and in-context examples used in our method can be found in the Appendix E.
4.1. Anchor Entity Identifying
@@
In our framework, the process of accurately identifying anchor entities from natural language questions is handled by the predictor agent. This process consists of the following two steps.
Candidate Entity Recognition. The predictor agent is associated with a LLM to recognize and extract keywords from the input question $q$, forming a keyword set $\mathcal{W}=\{w_{1},w_{2},\ldots,w_{n}\}$, where each $w_{i}$ denotes an entity mention relevant to $q$. As shown in Figure 2, given the question $q$: “What country bordering France contains an airport that serves Nijmegen?”, the predictor extracts $\mathcal{W}=\{\text{France},\text{Nijmegen}\}$. Considering the issues of entity abbreviations, aliases, long-tail entities, and potential spelling errors in open-world scenarios, directly using $\mathcal{W}$ for exact matching within a large-scale KG(Bollacker et al., 2008) often results in low recall and high computational cost. To handle this issue, AnchorRAG employs a robust entity linking mechanism based on exact semantic indexing. Specifically, AnchorRAG encodes all entity names in the KG with MiniLM(Wang et al., 2020), then conducts a FAISS(Douze et al., 2024) index from these embeddings, enabling efficient similarity search. On this basis, Predictor first cleans and proofreads the initial set $\mathcal{W}$ of query keywords, and filters out irrelevant interferences such as non-critical terms, and is prompted to correct potential spelling errors. Subsequently, with the inverted file and product quantization techniques provided by the FAISS library, Predictor can efficiently retrieves a set of the most relevant candidate entities $\tilde{\mathcal{E}}$ from the knowledge graph.
@@
Context-Aware Entity Disambiguation. The initial candidate set $\tilde{\mathcal{E}}$ often contains numerous entities with similar names that are irrelevant to the question’s context. Taking all of them into account will increase the overhead of subsequent retrieval and inevitably introduce noise. AnchorRAG is inspired by the assumption that we can effectively determine the entity’s relevance by measuring the semantic implied in its local graph structure. An entity is a more probable anchor if its neighboring relations are semantically aligned with the given question. Therefore, the Predictor employs a context-aware entity disambiguation strategy to precisely identify the most relevant candidates. Specifically, for each candidate entity $e\in\tilde{\mathcal{E}}$, the agent gathers its one-hop neighboring relations and obtains the set $\mathcal{R}(e)$. Then it invokes a pre-trained model SBERT (Reimers and Gurevych, 2019) to obtain vector representations for the question $q$ and each relation $r_{i}\in\mathcal{R}(e)$, denoted as $\mathbf{q}$ and $\mathbf{r}_{i}$ respectively. The relevance score $s(r_{i},q)$ between a relation and the question is then calculated using cosine similarity:
(2) $s(r_{i},q)=\frac{\mathbf{r}_{i}\cdot\mathbf{q}}{|\mathbf{r}_{i}||\mathbf{q}|},$
Furthermore, we define the final relevance score for the candidate entity $e$ by aggregating the scores of its most relevant neighboring relations:
(3) $\mathrm{Score}(e,q)=\frac{1}{k}\sum_{r\in\mathcal{N}(e)}s(r,q),$